Suppose that cost function is of a firm is given by C=Q3-4Q2+14Q+60. Then ,determine

a. Fixed Cost function and AFC at Q=2

b. TVC function and AVC at Q=2

c. MC function and MC at Q=2

d. Minimum average cost

1
Expert's answer
2021-12-20T10:18:11-0500

a. Fixed Cost function and AFC at Q=2

Fixed cost function; C=60

"AFC=\frac{Fixedcostfunction}{q}=\frac{60}{2}=30"


b. TVC function and AVC at Q=2

TVC=Q3-4Q2+14Q


"AVC=\frac{TVC}{Q}=\frac{Q^3-4Q^2+14Q}{Q}=Q^2-4Q+14"

"AVC=2^2-4\times2+14=18"


c. MC function and MC at Q=2

"MC=\frac{\Delta C}{\Delta Q}=3Q^2-8Q+14"


"MC=3(2)^2-8(2)+14=10"


d. Minimum average cost

"=\frac{\Delta TC}{\Delta Q}"


"=\frac{\Delta Q^3-4Q^2+14Q+60}{\Delta Q}"


"=Q^2-4Q+14+\frac{60}{Q}"